OpenStudy (anonymous):

What is the value of x, when x2 + 5 = 21?

OpenStudy (anonymous):

2x+5=21 -5 -5 2x=16 Divide both sides by 2 x=8

jimthompson5910 (jim_thompson5910):

Think she meant x^2 right?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

x^2 + 5 = 21 x^2 = 21 - 5 x^2 = 16 x = sqrt(16) or x = -sqrt(16) x = 4 or x = -4

jimthompson5910 (jim_thompson5910):

Don't forget about the plus minus since (-4)^2 = 16 as well as (4)^2 = 16